f29> Use the hypothetical data in the table to answer the questions. Snowfall (inches) City January February 36 78 Cincinnati, OH | 18 Boston, MA Buffalo, NY 97 38 What was the average snowfall in Cincinnati over this period? Round your answer to the first decimal. inches average snowfall: If the March data on snowfall is above the average that you calculated, what will happen to the average? 。The monthly averge will fall. O The monthly average will not change. O The monthly average will rise about us careers privacy policy terms of use contact us help
Answer #1

1.

The average snowfall in cincinnati over this period was  28 inches.

Explanation

In the city cincinnati, snowfall (inches) in january and february is 18 and 38 respectively :

\bg_white average \ snowfall= \frac{18+38}{2}= 28\ inches   

2.

If the march data on snowfall is above the average (i.e 28 inches), The monthly average will rise.

Explanation

If the march data on snowfall is above the average snowfall in january and february (i.e 28 inches), the monthly average will rise because, any snowfall data in march which is greater than 28 inches will lead to rise in monthly average.

For e.g

suppose, If we consider snowfall (inches) in march is 30 inches, than monthly average is :

Monthly \ average \ snowfall= \frac{18+38+31}{3}= 29\ inches

which is greater than 28 inches.

Therefore, If the march data on snowfall is above the average (i.e 28 inches), The monthly average will rise.